Compute for the pH and % ionization of a 0.75% v/v of a 150 ml diluted formic acid solution with a SG of 1.22 g/ml. kaformic acid = 1.77x10-4 

 

a. What is the weight of the formic acid?

b. What is the molecular weight of formic acid?

c. What is the equivalent weight of formic acid?

d. What is the normality of the dilute formic acid?

e. What is Hydrogen [H+] or hydronium [H3O+] concentration?

f. What is the pH of the diluted formic acid?

g. Whart is the % ionization constant of formic acid?


1
Expert's answer
2021-03-25T03:41:54-0400

Volume of formic acid = (0.75 / 100) × 150mL = 1.125mL

Weight of formic acid = SG × volume = 1.22 × 1.125

= 1.37g

Molecular weight of formic acid CH2O2 = 46g/mol

Equivalent weight = 46

Normality = 1.37g/( 1.125 × 10-3)L = 1218N